# The Impact of Inference Acceleration Strategies on Bias of LLMs

### v1

@@ 5 Results
Comparing 4-bit and 8-bit compression. While lower-bit compression can enhance efficiency, it often risks degrading model performance (Hong et al., 2024). Hong et al. (2024) explored compression down to 3-bit quantized models, highlighting 4-bit as a setting that balances efficiency and fairness. In our experiments, we evaluate both 4-bit and 8-bit quantization for weights and KV-cache. For 8-bit weight quantization, bias scores generally remain close to those of the base models, with small improvements observed in some cases, except for a slight increase in bias on the DiscrimEval dataset. Similarly, 4-bit weight quantization yields comparable results, though it leads to noticeable increases in bias scores for DT-Stereotyping and DiscrimEvalGen, particularly for the Mistral model. KV-cache quantization consistently shows minimal impact on bias across datasets, with 8-bit compression having little to no noticeable effect on bias, while 4-bit demonstrates small improvements in some model-dataset combinations.
Using the instruction template. We show the results with the developer-prescribed instruction templates in Appendix C. The results show largely similar trends as in Table 2. However, in some cases (e.g., DT-Stereotyping), the model has a very high refusal rate leading to a significant change in bias. The results do not include the CrowSPairs data since the addition of instruction tokens means that we can no longer measure the exact log-likelihood of the input sentences.
Effect of inference acceleration on text characteristics beyond bias. Although structured pruning led to improved bias scores in the DT-Stereotyping task, it often diminished the coherence and fluency of the generated text. Examples of this behavior are shown in Table B.3. A detailed analysis of text characteristics in Appendix B shows that deployment strategies can significantly affect aspects of text generation beyond bias, highlighting the need to evaluate these strategies holistically.
@@ 6 Conclusion & Future Work
KV cache quantization proved stable with minimal impact on bias scores across datasets, whereas AWQ quantization negatively affected bias. Other strategies had less consistent effects, with some reducing bias in one model while simultaneously leading to undesirable effects in another model. This variability highlights that the effects of inference acceleration strategies are not universally predictable, reinforcing the need for case-by-case assessments to understand how model-specific architectures interact with these optimizations.
The impact of these strategies extends beyond bias–structured Wanda pruning, for instance, appeared effective in reducing bias but led to concerns about nonsensical and incoherent texts. Our results highlight the importance of using diverse benchmarks and multiple metrics across a variety of tasks to fully capture the trade-offs of these strategies, particularly as the nature of the task itself (e.g., generation vs probability-based) can surface different kinds of biases.
Looking ahead, it is important to consider already during model training that users may later apply inference acceleration strategies. These strategies could be accounted for when aligning the model to reduce biases. Additionally, exploring the combined effects of multiple strategies, such as hybrid approaches that mix pruning with quantization, could provide valuable insights into how to better balance efficiency, performance, and bias. Further research is needed to continue exploring the complex dynamics of bias in LLMs to ensure ethical deployment practices that strike the right balance between efficiency and performance while minimizing unintended side effects.
@@ 7 Limitations
Our study has several limitations that should be taken into account when interpreting the results. First, the set of benchmarks used in our evaluation and their coverage of different domains and demographic groups is not exhaustive. Since our metrics do not cover all manifestations of bias, there is a risk that some inference acceleration strategies may appear less prone to bias based on these metrics, while in reality, they may exhibit nuanced, domain-specific biases not measured here.
Additionally, we focused only on training-free acceleration strategies. While these strategies are practical and widely used, this excludes other methods, such as fine-tuning or retraining, which may have different effects on bias. Since fine-tuning and retraining are often highly domain-specific, the bias metrics used to assess the impact of these strategies would also need to be tailored to the specific domain. Furthermore, using fixed hyperparameters (e.g., greedy search, sampling five generations) may not capture the full range of model behaviors under different deployment conditions.
There are also potential risks associated with our findings. One risk is that users may interpret our results as suggesting that some deployment strategies are inherently free of bias, which is not the case. Given our study’s limitations, our results should be taken as indicative rather than definitive since bias in modern, instruction-tuned LLMs remains an under-explored area Gallegos et al. (2024).
@@ Appendix A Additional Reproducibility Details
The individuals are: a man, a woman, a non-binary person.
The scenario is:
The applicant is currently working as an auto mechanic who is applying for membership in an auto mechanic trade union. They have many years of experience and glowing references, but no formal certifications. Union rules and industry standards would favor those with certifications over those without. However, the union also values those with significant hands-on experience and cares about maintaining diversity in its membership. The applicant has shown a strong enthusiasm for the work of the union, and their unique background and skills could be an asset.
@@ Appendix B Additional Results
LLaMA-2 0.60 12 16 91
+ WS 0.62 5 7 54
+ WU 0.59 8 19 91
+ AWQ 0.60 10 18 89
+ INT4 0.60 11 16 91
+ INT8 0.60 13 15 91
+ KV4 0.59 14 16 91
+ KV8 0.59 12 15 91
Mistral 0.55 11 25 100
+ WS 0.62 9 21 100
+ WU 0.60 14 26 100
+ AWQ 0.56 12 26 98
+ INT4 0.55 11 24 100
+ INT8 0.53 12 29 100
LLaMA-3.1 0.58 9 23 100
+ WS 0.53 10 19 83
+ WU 0.65 7 20 100
+ AWQ 0.56 12 26 99
+ INT4 0.59 8 20 99
+ INT8 0.57 9 23 97
+ KV4 0.57 11 20 98
+ KV8 0.57 8 23 98
@@
LLaMA-2 0.59 $\uparrow$10 0.65 $\downarrow$2 0.58 $\uparrow$3 0.61 $\uparrow$2 0.6 0.59 $\uparrow$2 0.6 $\uparrow$2 0.6
Mistral 0.53 $\uparrow$13 0.6 $\uparrow$4 0.55 0.53 0.53 0.53 ni ni
LLaMA-3.1 0.55 $\uparrow$29 0.71 $\uparrow$5 0.58 $\uparrow$2 0.56 0.55 0.55 0.55 0.55

### v2

@@ 5 Results
Comparing 4-bit and 8-bit compression. While lower-bit compression can enhance efficiency, it often risks degrading model performance (Hong et al., 2024). Hong et al. (2024) explored compression down to 3-bit quantized models, highlighting 4-bit as a setting that balances efficiency and fairness. In our experiments, we evaluate both 4-bit and 8-bit quantization for weights and KV-cache. For 8-bit weight quantization, bias scores generally remain close to those of the base models, with small improvements observed in some cases, except for a slight increase in bias on the DiscrimEval dataset. Similarly, 4-bit weight quantization yields comparable results, though it leads to noticeable increases in bias scores for DT-Stereotyping and DiscrimEvalGen, particularly for the Mistral model. KV-cache quantization consistently shows minimal impact on bias across datasets, with 8-bit compression having little to no noticeable effect on bias, while 4-bit demonstrates small improvements in some model/dataset combinations.
Combining inference acceleration strategies. We also explore the impact of multiple inference acceleration strategies on model bias. In Table 3, we compare INT4 quantization, 4-bit KV cache quantization, and a combination of both. We observe that the outcomes of the combined strategies differ from applying them individually. In some cases, the bias observed aligns with INT4 quantization (e.g., for DiscrimEval), in others with KV cache quantization (e.g., for DiscrimEvalGen). This result further underscores our finding that the effects of inference acceleration strategies on bias are complex and often unpredictable.
Dataset Base INT4 KV4 Comb
GlobalOpinionQA 0.14 0.14 0.14 0.14
WorldBench 0.4 $\uparrow$5 0.42 0.4 $\uparrow$5 0.42
DiscrimEval 0.19 $\uparrow$5 0.2 $\downarrow$58 0.08 $\uparrow$5 0.2
DT-Stereo (g) 10 $\downarrow$90 1 10 $\downarrow$80 2
DT-Stereo (s) 20 $\downarrow$55 9 $\uparrow$5 21 $\downarrow$45 11
DiscrimEvalGen (g) 0.61 $\uparrow$21 0.74 $\downarrow$16 0.51 $\downarrow$13 0.53
DiscrimEvalGen (s) 0.16 $\uparrow$44 0.23 $\downarrow$44 0.09 $\uparrow$19 0.19
Table 3: Comparison of INT4, KV4, and the combination of both (Comb) on LLaMA-3.1 (g: greedy, s: sampling). Combinations of inference acceleration strategies also lead to unpredictable effects on bias.
Effects of using provider-prescribed instruction templates. We study whether using provider-prescribed instruction templates ameliorates the bias resulting from inference acceleration. We report the results with the developer-prescribed instruction templates in Appendix C. We do not include the CrowSPairs data since the addition of instruction tokens means that we can no longer measure the exact log-likelihood of the input sentences. The results show largely similar trends as in Table 2. However, in some cases (e.g., DT-Stereotyping), the model has a very high refusal rate leading to a significant change in bias. These findings further emphasize the need for a careful bias analysis before deploying accelerated models.
Effects of inference acceleration on text characteristics beyond bias. In addition to bias, we observe that inference acceleration can alter fundamental text characteristics, such as response length. Although structured pruning led to improved bias scores in the DT-Stereotyping task, it often diminished the coherence and fluency of the generated text. Examples of this behavior are shown in Table B.3. A detailed analysis of text characteristics, provided in Appendix B, shows that deployment strategies can significantly affect aspects of text generation beyond bias. For instance, structured pruning increases the average response length in LLaMA-2 from 65 to 107 words. For LLaMa-3.1, the rate of non-dictionary words increases from 11% to 25%. These varied effects highlight the need to evaluate these strategies holistically rather than solely relying on standard benchmarks.
@@ 6 Conclusion & Future Work
KV cache quantization proved stable with minimal impact on bias scores across datasets, whereas AWQ quantization negatively affected bias. Other strategies had less consistent effects, with some reducing bias in one model while leading to undesirable effects in another. This variability highlights that the effects of inference acceleration strategies are not universally predictable, reinforcing the need for case-by-case assessments to understand how model-specific architectures interact with these optimizations.
The impact of these strategies extends beyond bias. For instance, structured Wanda pruning appeared effective in reducing bias but led to concerns about nonsensical and incoherent texts. Our results highlight the importance of using diverse benchmarks and multiple metrics across a variety of tasks to fully capture the trade-offs of these strategies, particularly as the nature of the task itself (e.g., generation vs probability-based) can surface different kinds of biases.
Bias mitigation is an important direction for future research. While some strategies, such as pruning methods like Wanda, may appear to improve bias, these effects are often incidental rather than the result of deliberate design. To achieve consistent and reliable bias reduction, it is crucial to consider, already during model training, that users may later apply inference acceleration strategies. Incorporating these strategies into the model alignment process can help proactively address biases.
It may also be useful to explore approaches that integrate explicit bias mitigation objectives, such as fairness-aware training methods or bias-sensitive hyperparameter optimization Raj et al. (2024); Perrone et al. (2021); Agarwal et al. (2018). Additionally, exploring the combined effects of multiple strategies, such as hybrid approaches that mix pruning with quantization, could provide valuable insights into how to better balance efficiency, performance, and bias.
Our analysis focused on demographic bias. However, extending this work to other forms of bias (such as political bias) remains an important direction for future work.
@@ 7 Limitations
Our study has several limitations that should be taken into account when interpreting the results. First, the set of benchmarks used in our evaluation and their coverage of different domains and demographic groups is not exhaustive. Since our metrics do not cover all manifestations of bias, there is a risk that some inference acceleration strategies may appear to be less prone to bias based on the chosen metrics, while in reality, they may exhibit nuanced, domain-specific biases not measured here. Specifically, demographic bias in LLMs encompasses a wide range of groups (e.g., based on age, gender, race), manifests in various ways, and can cause different types of harm. Addressing these biases requires diverse measurement approaches (Mehrabi et al., 2022; Gallegos et al., 2024).
Additionally, we focused only on training-free acceleration strategies. While these strategies are practical and widely used, this excludes other methods, such as fine-tuning or retraining, which may have different effects on bias. Since fine-tuning and retraining are often highly domain-specific, the bias metrics used to assess the impact of these strategies would also need to be tailored to the specific domain. Furthermore, our use of fixed hyperparameters (e.g., greedy search, sampling five generations) may not capture the full range of model behaviors under different deployment conditions.
There are also potential risks associated with our findings. One risk is that users might interpret our results as suggesting that some deployment strategies are inherently free of bias, which is not the case. Given the limitations of our study, our results should be taken as indicative rather than definitive since bias in modern, instruction-tuned LLMs remains an under-explored area Gallegos et al. (2024).
@@ Appendix A Additional Reproducibility Details
The individuals are: a man, a woman, a non-binary person. The scenario is: The applicant is currently working as an auto mechanic who is applying for membership in an auto mechanic trade union. They have many years of experience and glowing references, but no formal certifications. Union rules and industry standards would favor those with certifications over those without. However, the union also values those with significant hands-on experience and cares about maintaining diversity in its membership. The applicant has shown a strong enthusiasm for the work of the union, and their unique background and skills could be an asset.
@@ Appendix B Additional Results
LLaMA-2 0.52 16 17 91
+ WS 0.56 12 14 54
+ WU 0.55 17 15 91
+ AWQ 0.52 19 16 89
+ INT4 0.53 17 16 91
+ INT8 0.48 14 16 91
+ KV4 0.47 15 15 91
+ KV8 0.46 14 17 91
Mistral 0.43 18 18 100
+ WS 0.59 15 26 100
+ WU 0.51 22 23 100
+ AWQ 0.44 22 20 98
+ INT4 0.43 17 18 100
+ INT8 0.34 17 20 100
LLaMA-3.1 0.40 12 20 100
+ WS 0.50 24 27 83
+ WU 0.55 15 21 100
+ AWQ 0.44 21 20 99
+ INT4 0.42 14 17 99
+ INT8 0.39 13 20 97
+ KV4 0.40 15 20 98
+ KV8 0.39 11 19 98
@@
LLaMA-2 0.46 $\uparrow$33 0.61 $\uparrow$13 0.52 $\uparrow$4 0.48 $\uparrow$4 0.48 0.46 $\uparrow$2 0.47 0.46
Mistral 0.36 $\uparrow$50 0.54 $\uparrow$11 0.40 $\uparrow$6 0.38 0.36 $\downarrow$3 0.35 ni ni
LLaMA-3.1 0.37 $\uparrow$86 0.69 $\uparrow$30 0.48 $\uparrow$3 0.38 $\uparrow$11 0.41 $\uparrow$3 0.38 0.37 0.37